OpenStudy (anonymous):

Solve –4(x + 4) = 16 i forgot how to do the step can you please help me:)

OpenStudy (anonymous):

if four times a number is 16 then the number is 16 divided by 4

OpenStudy (anonymous):

don't i have to do distributive property ?

hartnn (hartnn):

its easier to divide the whole equation by -4 first, then u don't need to use distributive property ...

OpenStudy (anonymous):

x=-8 right

OpenStudy (anonymous):

so i divide negative four by 16? or the opposite ?

hartnn (hartnn):

u divide 16 by (-4)

OpenStudy (anonymous):

i get -4?

hartnn (hartnn):

yes :) so u have x+4=-4 can u solve further ?

OpenStudy (anonymous):

-4(-8+4)=16 -4(-4)=16 16=16

OpenStudy (anonymous):

i subtracted both sides by 4 and got -8 for my final answer ; x= -8

hartnn (hartnn):

thats correct :) now u understand ? ask if still in any doubts anywhere.....

OpenStudy (anonymous):

Yayy:) well one little question, i didn't really get why i had to divide 16 by 4?

OpenStudy (anonymous):

you dont

OpenStudy (anonymous):

i got alittle confused

OpenStudy (anonymous):

to be honest

OpenStudy (anonymous):

oh actually you do

hartnn (hartnn):

since the main aim was to isolate x and -4 was "multiplied" by (x+4), so to isolate, u divide both sides by -4 so that only x+4 remain on left. got this ?

OpenStudy (anonymous):

-4(x+4)=16 x+4=-4 x+4-4=-4-4 x=-8

OpenStudy (anonymous):

@hartnn yeaah i understand now:) @BlizzardofFire thanks!